# Training Verification-Friendly Neural Networks via Neuron Behavior Consistency

### v1

@@ Introduction
Table 2 shows the results of networks trained with our method combined with other methods on MNIST dataset. All examples show that our method combined with existing methods improves the ratio of stable neurons. In most cases, our method combined with existing methods increases the ratio of verified properties (UNSAT%) and reduces the time required for verification greatly.
It is worth noting that when the baseline method is combined with our method, the performance is particularly outstanding in larger network structures. For example, in the M3 model. Our method combined with Madry and RS methods respectively increased the verified ratio by 83.2% and 87.0% at $\varepsilon=0.2$ and also decreased the time by 94.4 seconds and 102.5 seconds. This means that our method makes it possible to verify larger network structures. The same results are also reflected in the Fashion-MNIST and CIFAR-10 datasets provided in the supplementary material.
Combining our method with existing methods sacrifices some accuracy and adversarial accuracy, but improves verification and neural stability. In particular, the improvement is more pronounced in larger models. This means that our method makes it possible to verify larger network structures using existing verification tools. The same results are also reflected in the Fashion-MNIST and CIFAR datasets provided in the supplementary material.
@@
Answer RQ2: Our method combined with other methods demonstrates to improve the verification-friendly properties of the network, however, a trade-off between accuracy and verifiability is required.
@@
Figure 3: Networks trained with $\varepsilon={2}/{255}$ on CIFAR-10 dataset and evaluated on $\varepsilon={2}/{255}$. Methods unable to achieve given accuracy range are omitted. ’Natural’ means the network are trained with only cross-entropy loss.
The accuracy of networks trained using default parameters varied significantly on the CIFAR-10 dataset. To compare the performance of networks with close accuracy, we fine-tuned the parameters to ensure that the accuracy of each method fluctuated within a certain range.
Figure 3 shows the results of networks trained on CIFAR-10. Larger perturbation radii are very challenging for current verification methods, so we choose a smaller perturbation radius to train and evaluate the networks.
In terms of UNSAT%, our method performs best in different network models and different accuracy ranges and generally requires less verification time for timeouts and UNSAT results than other methods, indicating that our networks are easier to verify. When the models reach higher accuracy, our method has a slightly longer average time, as other methods are more easily to find counterexamples using attack methods, while our method is less vulnerable to attacks.
Our accuracy under PGD is generally higher than other methods, with a few exceptions, where it is slightly lower than the best-performing method. Our method consistently maintains a higher ratio of stable neurons than other methods, especially in the C3 model, our method maintains a high ratio of stable neurons of up to 80.1% while maintaining a high verification accuracy, indicating that the upper bound of the search space is significantly reduced.
@@
Answer RQ3: Experiments show that our method is more verification-friendly than other methods at higher and close accuracy levels, as our method consistently maintains a high ratio of stable neurons, and relative robustness and requires less verification time.
@@
Traditional adversarial methods such as TRADES (Zhang et al. 2019b) and Madry et al. (Madry et al. 2018) use adversarial training to improve the robustness of neural networks. However, these methods most impose constraints on the network output, while our method imposes neuron behavior constraints on all layers of the network. Recent works, HYDRA (Sehwag et al. 2020) integrates the pruning process with adversarial training, using a criterion that considers adversarial robustness during the pruning decision, leading to more compact yet robust models and Wu et al. (Wu, Xia, and Wang 2020) found that adversarial weight perturbation can improve the robustness of the network.
Certified training (De Palma et al. 2022; Mirman, Gehr, and Vechev 2018; Jovanovic et al. 2022; Zhang et al. 2019a; Xu et al. 2020) introduces interval bound propagation (IBP) into the training process to improve the robustness of the network but suffers from the long training time.
The ReLU Stable method (Xiao et al. 2019) ensures that the upper and lower bounds of the neurons have the same sign, pushing the upper and lower bounds away from zero in the same direction. This method can also incorporate ternary loss to train a SAT-friendly network (Narodytska et al. 2019). In contrast, our method emphasizes the stability of neurons before and after perturbation, focusing on the similarity and consistency of neuron behavior. Furthermore, RS loss requires additional calculations for each neuron’s bounds, whereas our method does not require extra information, making it easier to implement. Linearity grafting (Chen et al. 2022c) replaces unstable neurons with linear neurons to improve the robustness of the network. However, this method will modify network architectures, while our method keeps the architecture unchanged. MILP-based method (Baninajjar, Rezine, and Aminifar 2023) uses MILP to post-process the network to make it sparse to improve the verification efficiency, which may be time-consuming. Pruning-based methods (Xiao et al. 2019; Xu et al. 2024) heuristically prune the neurons that are inactive or unstable and have little impact on the network’s performance. Bias shaping method (Xu et al. 2024) changing the bias of the unstable neurons during training to stabilize them. These methods are either post-processing methods or training tricks that can be incorporated with other methods to further improve the verifiability of networks.
@@ Conclusion
In this work, we propose a novel training method to train a verification-friendly neural network by preserving the neuron behavior consistency of the network. Our experiments on the MNIST, Fashion-MNIST, and CIFAR-10 datasets demonstrate that our method consistently enhances the network’s verification-friendliness, as evidenced by a stable neuron ratio, comparable robustness, and faster verification speed across different perturbation radii. Additionally, existing methods, when combined with ours, show improved verification efficiency. Our method also speeds up the verification process while maintaining the accuracy of the model, which is not commonly achieved by existing methods.

### v2

@@ Introduction
Table 2 shows the results of networks trained with our method combined with other methods on MNIST dataset. All the results show that our method combined with existing methods improves the ratio of stable neurons. In most cases, our method combined with existing methods increases the ratio of verified properties (UNSAT%) and reduces the time required for verification greatly.
It is worth noting that when our method is combined with a baseline method, the performance is particularly outstanding in larger network structures. For example, in the M3 model. Our method combined with the Madry and RS methods respectively increased the verified ratio by 83.2% and 87.0% at $\varepsilon=0.2$, while also reducing verification time by 94.4 seconds and 102.5 seconds. This means that our method makes it possible to verify larger network structures. The same results are also reflected in the Fashion-MNIST and CIFAR-10 datasets provided in the supplementary material.
Combining our method with existing methods sacrifices some accuracy and adversarial accuracy but leads to improved verification and neural stability. The improvement is particularly significant for larger models, meaning that our method enables the verification of larger network structures using existing verification tools. Similar results are also observed in the Fashion-MNIST and CIFAR-10 datasets, as shown in the supplementary material.
@@
Answer to RQ2: Our method, when combined with other methods, improves the verification-friendly properties of the network. However, a trade-off between accuracy and verifiability is required.
@@
Figure 3: Networks trained with $\varepsilon={2}/{255}$ on CIFAR-10 dataset and evaluated on $\varepsilon={2}/{255}$. Methods unable to achieve a given accuracy range are omitted. ’Natural’ means the network is trained with only cross-entropy loss.
The accuracy of networks trained with default parameters varied significantly on the CIFAR-10 dataset. To compare the performance of networks with similar accuracy, we fine-tuned the parameters to ensure that the accuracy of each method remained within a certain range.
Figure 3 shows the results of networks trained on CIFAR-10. Larger perturbation radii are very challenging for current verification methods, so we chose a smaller perturbation radius to train and evaluate the networks.
For UNSAT%, our method performs best across different network models and different accuracy ranges and generally requires less verification time for timeouts and UNSAT results than other methods, indicating that our networks are easier to verify. When the models reach higher accuracy, our method has a slightly longer average time, as other methods are more easily to find counterexamples using attack methods, while our method is less vulnerable to attacks.
Our accuracy under PGD is generally higher than that of other methods, with a few exceptions where it is slightly lower than the best-performing method. Our method consistently maintains a higher ratio of stable neurons than other methods. Notably, in the C3 model, our method maintains a high stable neuron ratio of up to 80.1% while preserving high verification accuracy. This indicates that the upper bound of the search space is significantly reduced.
@@
Answer to RQ3: Experiments show that our method is more verification-friendly than other methods at higher and close accuracy levels, as it consistently maintains a high ratio of stable neurons, greater robustness, and requires less verification time.
@@
Traditional adversarial methods such as TRADES (Zhang et al. 2019b) and Madry (Madry et al. 2018) use adversarial training to improve the robustness of neural networks. However, these methods primarily impose constraints on the network’s output, while our method imposes neuron behavior constraints on all the network layers. Recent works, such as HYDRA(Sehwag et al. 2020), integrate the pruning process with adversarial training, using a criterion that considers adversarial robustness during the pruning decision. This leads to more compact yet robust models. Wu et al. (Wu, Xia, and Wang 2020) also found that perturbing the weight can improve the robustness of the network.
Certified training (De Palma et al. 2022; Mirman, Gehr, and Vechev 2018; Jovanovic et al. 2022; Zhang et al. 2019a; Xu et al. 2020) introduces interval bound propagation (IBP) into the training process to improve network robustness but suffers from long training times.
The ReLU Stable method (Xiao et al. 2019) ensures that the upper and lower bounds of the neurons have the same sign, pushing them away from zero in the same direction. This method can also incorporate ternary loss to train a SAT-friendly network (Narodytska et al. 2019). In contrast, our method emphasizes the stability of neurons before and after perturbation, focusing on the consistency of neuron behavior. Furthermore, RS loss requires additional calculations for each neuron’s bounds, whereas our method does not require extra information, making it easier to implement.
Linearity grafting (Chen et al. 2022c) replaces unstable neurons with linear neurons to improve the robustness of the network. However, this method modifies the network architecture, whereas our method preserves the architecture.
The MILP-based method (Baninajjar, Rezine, and Aminifar 2024) uses MILP to post-process the network and make it sparse, improving verification efficiency. However, this may be time-consuming.
Pruning-based methods (Xiao et al. 2019; Xu et al. 2024) heuristically prune inactive or unstable neurons with little impact on the network’s performance. The bias shaping method (Xu et al. 2024) changes the bias of unstable neurons during training to stabilize them. These methods are either post-processing techniques or training tricks that can be incorporated with other methods to improve the verifiability of networks further.
@@ Conclusion
In this work, we propose a novel training method to develop a verification-friendly neural network by preserving neuron behavior consistency. Our experiments on the MNIST, Fashion-MNIST, and CIFAR-10 datasets demonstrate that our method consistently enhances the network’s verification-friendliness, as evidenced by a higher stable neuron ratio, comparable robustness, and faster verification speed across different perturbation radii. Additionally, when combined with existing methods, our method shows improved verification efficiency. Our method also accelerates the verification process while maintaining model accuracy, which is a result that is rarely achieved by existing methods. In future work, we plan to explore the application of our method to more complex network architectures and datasets, as well as explain the underlying mathematical machinery of our method in more detail.
@@ Acknowledgements
Appendix A Detailed Experimental Setup
Dataset Selection.
We select three datasets: Fashion-MNIST, MNIST, and CIFAR-10 datasets for our experiments. The MNIST and CIFAR-10 datasets are widely used in early adversarial training research (Madry et al. 2018; Zhang et al. 2019b) and similar studies (Xiao et al. 2019; Xu et al. 2024). To demonstrate the generality of our method, we additionally included the Fashion-MNIST dataset. Fashion-MNIST is also a dataset for a ten-class classification task like the MNIST dataset, but the images in the Fashion-MNIST dataset are more complex and challenging compared to the MNIST dataset.
Network Architecture Selection.
We select the following network architectures based on the network sizes used in previous works (Xiao et al. 2019; Xu et al. 2024), the verification capabilities of neural network verification tools, and the generality of the networks. M1 and C1 are smaller networks, M2 and C2 are medium-sized networks, and M3 and C3 are larger networks. Compared to previous works, we mostly use convolutional networks instead of fully connected networks. This is because the use of convolutional networks is more common in real applications. Moreover, M3 and C3 are even larger than the largest network structures used in similar works (Xiao et al. 2019; Xu et al. 2024) and are very close to the limits of neural network verification tools. Table 3 provides detailed descriptions of the network architectures used in our experiments.
Training Details.
For the MNIST and the Fashion-MNIST datasets, we directly use the original pictures to train each network. For the CIFAR-10 dataset, we randomly crop images to 32x32 pixels and augment the data by flipping images horizontally with a probability of 0.5. We also add padding of 4 pixels on each side before cropping.
We use the Adam optimizer to train all networks with a batch size of 128. The learning rate is set to 0.0001 for the MNIST and Fashion-MNIST datasets and 0.00001 for the CIFAR-10 dataset. We use the same number of training iterations for all datasets, which is 400 iterations. The random seed is set to 0 for reproducibility. All PGD-like adversarial training process uses the same hyperparameters as 10 steps and a step size of $\varepsilon/10$.
In the last epoch experiment, for the RS (Xiao et al. 2019), Madry (Madry et al. 2018), and TRADES (Zhang et al. 2019b) methods, we used the same hyperparameters as in the original papers. As for our model, we use the hyperparameters $\beta=1$ for M1, M2 and M3 models, and $\beta=2,5,3$ for C1, C2, and C3 models, respectively. We select the hyperparameters based on the performance of the model on the validation set (generated by the original training set).
In the combination experiment, we find that different methods have different performances when using different combination orders. Based on the performance of the model on the validation set. For the RS method, we first train 200 epochs using the original RS loss and then combine it with our NBC loss for another 200 epochs. For the Madry method, we use the method combined with NBC loss for 200 epochs and then use the original Madry loss for another 200 epochs. For the TRADES method, we use the method combined with the NBC loss to train directly for 400 epochs. The parameters are the same as the original training process.
The application of these methods results mainly in a decrease in accuracy compared to direct training without these methods, which may be unacceptable in practical applications. Therefore, our objective is to evaluate the effectiveness of these methods while maintaining relatively high accuracy. We specified two sets of accuracy ranges for each model: one where the accuracy difference from the ’Natural’ method is within 1%-2%, and another where the difference is within 6%-7%. We believe this setup effectively evaluates the effectiveness of each method under different accuracy requirements. Besides, as shown in Table 1 and 5, various metrics drop as the robustness radius increases, making the differences in metrics between various methods less apparent. The use of $\varepsilon=2/255$ to train the model and the same $\varepsilon$ to verify the model reveals the differences more clearly and better illustrates the issue. In this experiment, we initially trained a baseline network using only cross-entropy loss (referred to as ’Natural’ in the tables and figures). We then used this model as a pre-trained baseline and applied various methods to fine-tune it separately, aiming to achieve a relatively high accuracy.
Verification Tasks.
To demonstrate that our methods have generalization capabilities, we train on the training set and generate local robustness verification tasks under different perturbation radii on the test set. We use the same verification tasks for all methods to ensure fairness.
Dataset Architecture Name #Parameters Network Architecture
(Fashion-) MNIST Cnn 4layer (M1) 166,406 Conv2d(1, 16, (4, 4), (2, 2), (1, 1))
Conv2d(16, 32, (4, 4), (2, 2), (1, 1))
Linear(1568, 100), Linear(100, 10)
Relu Stable (M2) 171,158 Conv2d(1, 16, (5, 5), (2, 2), (2, 2))
Conv2d(16, 32, (5, 5), (2, 2), (2, 2))
Linear(1568, 100), Linear(100, 10)
Conv Big (M3) 1,974,762 Conv2d(1, 32, (3, 3), (1, 1), (1, 1))
Conv2d(32, 32, (4, 4), (2, 2), (1, 1))
Conv2d(32, 64, (3, 3), (1, 1), (1, 1))
Conv2d(64, 64, (3, 3), (2, 2), (1, 1))
Linear(3316, 512), Linear(512, 512)
Linear(512, 10)
CIFAR-10 Marabou medium (C1) 165,498 Conv2d(3, 16, (4, 4), (2, 2), (2, 2))
Conv2d(16, 32, (4, 4), (2, 2), (2, 2))
Linear(1152, 128), Linear(128, 64)
Linear(64, 10)
Marabou large (C2) 338,346 Conv2d(3, 16, (4, 4), (2, 2), (2, 2))
Conv2d(16, 32, (4, 4), (2, 2), (2, 2))
Linear(2304, 128), Linear(128, 64)
Linear(64, 10)
Conv Big (C3) 2,466,858 Conv2d(3, 32, (3, 3), (1, 1), (1, 1))
Conv2d(32, 32, (4, 4), (2, 2), (1, 1))
Conv2d(32, 64, (3, 3), (1, 1), (1, 1))
Conv2d(64, 64, (3, 3), (2, 2), (1, 1))
Linear(4096, 512), Linear(512, 512)
Linear(512, 10)
Table 3: Network Architectures for Experiments
Model M1 M2 M3
Method T M R O T M R O T M R O
Test Acc. 84.0 84.4 81.3 82.1 83.8 86.0 80.5 78.3 87.2 89.3 86.2 81.6
$\varepsilon$=0.1 UNSAT53.5 30.6 70.1 69.1 49.3 2.0 70.0 65.5 16.3 0.2 27.1 63.3
Stable36.6 66.7 68.0 78.9 31.1 47.7 65.2 86.2 4.4 3.8 6.7 84.4
Time (s) 25.2 26.5 9.5 6.8 29.7 19.3 9.4 8.0 79.7 33.0 68.4 20.4
[…]
$\text{PGD}_{100}$ 70.5 49.5 73.8 71.8 68.9 30.5 73.6 71.1 77.5 31.8 78.1 74.7
$\varepsilon$=0.2 UNSAT24.8 6.0 55.4 54.5 9.9 0.1 54.2 46.0 0.0 0.0 0.8 42.4
Stable15.1 34.0 37.7 63.3 9.8 17.3 34.9 70.1 0.0 0.2 0.1 63.2
Time (s) 48.6 42.7 15.7 9.5 61.2 13.8 16.9 22.0 82.0 32.1 83.6 36.9
[…]
$\text{PGD}_{100}$ 62.7 42.5 66.3 61.8 62.6 32.5 67.1 65.0 68.8 31.2 70.8 67.9
$\varepsilon$=0.3 UNSAT5.3 0.4 31.2 35.6 0.2 0.0 30.0 17.4 0.0 0.0 0.0 17.4
Stable6.2 15.2 14.4 54.2 1.9 5.8 12.6 57.4 0.0 0.0 0.0 43.8
Time (s) 52.6 34.6 30.6 16.4 49.5 10.4 33.2 41.2 62.8 34.6 69.2 51.0
[…]
$\text{PGD}_{100}$ 53.1 36.7 57.3 49.4 51.1 31.7 58.9 59.5 58.3 36.8 62.2 59.5
Table 4: Networks trained with $\varepsilon=0.3$ on Fashion-MNIST datasets. The best results are highlighted in bold. Verified under $\varepsilon$=0.1, 0.2, 0.3. T:TRADES, M:Madry, R:ReLU Stable, O:Ours.
Model C1 C2 C3
Method T M R O T M R O T M R O
Test Acc. 59.2 62.8 56.0 51.8 60.9 69.5 61.8 54.9 69.5 75.5 74.1 64.9
$\frac{2}{255}$ UNSAT46.5 49.0 48.5 42.0 50.0 51.0 47.5 44.0 24.0 9.0 12.0 39.0
Stable79.3 76.2 80.4 91.6 77.6 69.6 74.8 94.2 53.4 42.3 44.1 80.3
Time (s) 13.8 14.8 11.1 9.8 17.1 21.2 24.7 14.2 77.7 112.9 105.2 46.8
[…]
$\text{PGD}_{100}$ 51.3 53.8 49.9 50.9 53.5 59.3 55.1 48.2 62.7 64.5 65.3 57.8
$\frac{4}{255}$ UNSAT22.5 17.0 30.0 31.0 17.5 8.0 14.5 28.0 1.0 0.5 1.0 2.5
Stable48.9 40.0 50.7 80.5 40.4 24.1 33.6 84.9 7.3 2.8 3.3 50.8
Time (s) 39.3 56.3 33.1 17.9 61.5 80.4 61.4 27.3 101.8 104.3 104.7 90.4
[…]
$\text{PGD}_{100}$ 43.6 44.5 44.0 41.8 46.6 48.5 48.3 41.5 54.8 52.8 55.6 50.8
$\frac{6}{255}$ UNSAT5.0 2.5 9.0 17.0 1.0 1.0 1.5 12.5 1.0 0.0 0.0 1.0
Stable23.4 15.9 24.3 68.0 14.0 5.7 9.7 73.7 1.1 0.4 0.4 44.3
Time (s) 49.0 56.5 54.9 33.4 69.3 62.6 61.7 39.9 89.6 85.4 82.4 78.6
[…]
$\text{PGD}_{100}$ 36.8 35.5 38.5 33.0 39.5 37.9 41.1 36.0 47.4 40.9 45.4 44.4
$\frac{8}{255}$ UNSAT1.0 0.5 1.0 7.5 0.5 0.5 0.5 2.5 0.0 0.0 0.0 0.5
Stable10.1 5.8 10.8 57.7 4.2 1.2 2.5 65.3 0.2 0.1 0.0 43.3
Time (s) 45.8 44.0 54.3 34.3 54.4 43.4 53.3 44.9 75.0 62.7 65.2 66.5
[…]
$\text{PGD}_{100}$ 29.9 27.4 32.9 25.3 32.7 28.3 33.9 30.2 39.8 30.5 35.3 38.0
Table 5: Networks trained with $\varepsilon=8/255$ on CIFAR-10 dataset with default parameters. The best results are highlighted in bold. Verified under $\varepsilon\in\{2/255,4/255,6/255,8/255\}$. T: TRADES, M: Madry, R: ReLU stable, O: Our method.
Model M1 M2 M3
Method T* M* R* T* M* R* T* M* R*
Test Acc. -7.7% -8.8% -2.9% -3.6% -10.0% -3.7% -5.7% -13.0% -7.5%
$\varepsilon$=0.1 UNSAT+1.9 +6.0 -8.3 -2.6 +28.0 -6.4 +19.5 +48.5 +37.6
Stable+52.7 +17.2 +14.8 +53.2 +28.9 +21.4 +71.6 +73.9 +82.6
Time (s) -17.6 -5.0 -3.6 -14.4 +8.5 -4.2 -32.7 -5.2 -57.1
[…]
$\text{PGD}_{100}$11.1 +7.6 -9.5 -11.4 +27.6 -7.9 -9.6 +36.5 -10.7
$\varepsilon$=0.2 UNSAT+11.9 -4.5 -7.7 -3.4 +0.5 -2.7 +0.1 +0.5 +49.1
Stable+62.1 +29.5 +28.8 +54.0 +42.7 +38.5 +47.2 +45.8 +76.4
Time (s) -34.5 -19.1 -8.0 -20.7 +14.4 -9.9 -11.4 +37.6 -66.6
[…]
$\text{PGD}_{100}$16.8 -8.1 -14.7 -21.1 +2.7 -12.0 -11.0 +29.6 -13.8
$\varepsilon$=0.3 UNSAT+11.7 -0.4 -2.4 -0.1 0.0 +5.5 0.0 0.0 +30.6
Stable+58.0 +40.0 +36.5 +49.0 +48.6 +47.8 +17.8 +2.1 +63.6
Time (s) -30.2 -25.6 -17.4 -29.7 -1.6 -22.5 -11.2 -16.5 -43.8
[…]
$\text{PGD}_{100}$19.5 -32.4 -19.5 -31.5 -26.6 -15.6 -14.5 +13.8 -17.2
Table 6: Networks trained with each method combined with our method on the Fashion-MNIST dataset with $\varepsilon=0.3$. Improved results are highlighted in bold. T*: TRADES+Ours, M*: Madry+Ours, R*: ReLU Stable+Ours.
Model C1 C2 C3
Method T* M* R* T* M* R* T* M* R*
Test Acc. -9.2% -14.2% -8.1% +5.4% -11.0% -0.1% -0.2% -9.3% -13.4%
$\varepsilon$=$\frac{2}{255}$ UNSAT-6.0 -7.5 -7.5 -9.5 -4.0 -2.5 -0.9 +35.5 +32.5
Stable+17.2 +17.8 +13.5 +14.3 +22.2 +17.2 +9.5 +45.7 +47.3
Time (s) -3.9 -6.3 -3.5 +8.2 -6.7 -11.0 +3.5 -85.4 -84.1
[…]
$\text{PGD}_{100}$7.9 -10.3 -6.5 -1.5 -7.8 -6.3 -0.5 -4.8 -18.6
$\varepsilon$=$\frac{4}{255}$ UNSAT+7.5 +15.0 +5.7 -10.5 +17.5 +2.5 +0.0 +2.5 +5.0
Stable+42.7 +45.7 +35.3 +36.6 +52.4 +46.5 +15.2 +58.1 +76.1
Time (s) -23.3 -38.5 -20.6 -14.1 -44.9 -31.6 +1.0 -18.0 -46.5
[…]
$\text{PGD}_{100}$6.6 -5.9 -5.1 -9.2 -3.9 -12.0 -0.5 +0.3 -13.9
$\varepsilon$=$\frac{6}{255}$ UNSAT+13.5 +21.0 +15.5 0.0 +3.5 +0.5 -1.0 +0.5 +10.6
Stable+62.2 +59.6 +53.0 +47.7 +50.1 +56.0 +16.5 +51.6 +77.3
Time (s) -22.8 -36.2 -35.4 -38.4 -5.2 -27.3 -2.0 -9.1 -33.0
[…]
$\text{PGD}_{100}$5.4 -1.5 -4.2 -14.7 -0.4 -15.9 -0.6 +4.9 -8.8
$\varepsilon$=$\frac{8}{255}$ UNSAT+9.5 +10.1 +13.0 -0.5 +0.5 -0.5 0.0 0.0 +2.0
Stable+67.9 +57.6 +56.7 +50.8 +39.1 +52.2 +16.1 +34.0 +69.2
Time (s) -16.6 -7.2 -22.4 -33.3 +0.9 -31.3 -7.5 +3.5 -13.9
[…]
$\text{PGD}_{100}$3.6 +2.0 -2.7 -17.7 +2.2 -17.8 -0.1 +7.9 -3.2
Table 7: Networks trained with each method combined with our method on the CIFAR-10 dataset with $\varepsilon=8/255$. Improved results are highlighted in bold. T*: TRADES+Ours, M*: Madry+Ours, R*: ReLU Stable+Ours.
[…]
C1 TRADES 69.8 38.5 77 43.4 67.3 55.8 55.8 53.0
Madry 69.6 39.5 79 41.7 63.8 54.3 54.3 50.8
RS 69.3 44.5 89 30.9 44.4 56.7 56.7 60.9
Ours 69.0 46.0 92 27.3 37.8 56.4 56.4 77.6
RS 75.0 1.5 3 49.9 171.9 26.4 26.4 8.5
Ours 75.0 33.0 66 54.4 87.0 56.1 56.1 46.3
Natural 77.6 0.5 1 49.6 178.4 10.0 10.0 5.3
C2 TRADES 74.7 35.0 70 55.0 85.3 60.4 60.4 46.9
Madry 74.4 28.5 57 65.5 103.1 58.7 58.8 39.4
RS 74.4 41.5 83 53.9 77.1 61.4 61.4 49.4
Ours 74.0 45.5 91 38.8 55.0 61.3 61.3 56.8
RS 79.3 1.0 2 54.1 173.9 28.5 28.4 3.7
Ours 79.3 12.0 24 87.9 147.3 57.3 57.2 26.3
Natural 80.6 0.5 1 53.3 181.0 10.0 10.0 2.8
C3 TRADES 77.6 5.5 11 112.8 165.9 64.0 64.0 27.7
Madry 77.8 1.0 2 113.9 177.2 60.1 60.1 17.0
RS 78.0 2.5 5 122.2 176.8 61.1 61.2 21.9
Ours 77.0 25.6 51 95.4 126.2 66.7 66.7 80.1
RS 83.7 0.0 0 65.5 180.4 30.3 30.3 0.1
Ours 83.3 1.0 2 115.6 177.3 61.0 61.0 5.6
Natural 84.2 0.0 0 68.0 184.1 10.0 10.0 0.0
Table 8: Networks trained on CIFAR-10 dataset with $\varepsilon=2/255$. The best results are highlighted in bold. ’Natural’ denotes the accuracy with only CE loss used.
Appendix B Detailed Experimental Results
The following tables provide detailed results of the experiments conducted for RQ1, RQ2, and RQ3. Typically, Table 4, 5 and Table 6, 7 show the results of networks trained for RQ1 and RQ2, on Fashion-MNIST and CIFAR-10 datasets, respectively. Table 8 shows the results of networks trained for RQ3 on CIFAR-10 datasets.
Extra Results for RQ1
Table 4 shows the evaluation results of networks trained in Fashion-MNIST.
In the M1 model, our method demonstrates moderate accuracy. Although our method is not the best in terms of UNSAT%, it is within 1% of the best performing RS method. Our method performs best in terms of UNSATat $\varepsilon=0.3$. Our method consistently maintains the best performance in terms of stability, with a stable ratio of over 50%. Our method is the fastest in terms of verification time, and its advantage becomes more pronounced as the perturbation radius increases. Our method performs moderately under the PGD attack.
In the M2 model, our method still performs best in terms of Stable%. However, various methods have their advantages and disadvantages.
It should be noted that in the M3 model, our method performs best in terms of UNSAT%, Stable%, and the average time of UNSAT and timeout issues. Especially at $\varepsilon=0.3$, our method performs best in terms of UNSAT%, reaching 17.4%, while networks trained with other methods cannot verify any UNSAT properties. Our method performs moderately under the PGD attack.
Table 5 shows the evaluation results of networks trained on CIFAR-10. Although our method does not perform well on test set accuracy compared to other approaches, it outperforms other metrics.
Our method demonstrates superior performance in terms of the average verification speed and the average speed of UNSAT and Timeout issues, requiring less verification time on almost all verification tasks. This is especially evident in the C1 and C2 models, where our method consistently outperforms other approaches in terms of verification time. Although the accuracy under 100 steps of the PGD attack is not the best-performing method, our method still achieves competitive PGD accuracy.
Our method outperforms others in terms of verification accuracy as adversarial perturbations increase, affirming its ability to maintain verification-friendly properties. Our approach consistently maintains a high ratio of stable neurons, even under the most challenging perturbation radii. This is particularly evident in the C2 and C3 models, where our method achieves a stable neuron ratio of over 50% at the highest perturbation radius. In contrast, other methods almost lose all stable neurons at the same perturbation radius.
Extra Results for RQ2
Table 6 and Table 7 show the results of networks trained with each method combined with our method on the Fashion-MNIST and CIFAR-10 datasets, respectively.
Our method significantly increases the ratio of stable neurons, which implies a smaller theoretical upper bound for solving. In most cases, the verification time is greatly reduced after combining our method. The UNSATslightly decreases when combined with our method at the smallest $\varepsilon$, but as the robustness radius $\varepsilon$ increases, our method performs almost the best in terms of UNSAT%. It is worth noting that, in the largest model, combining our method almost improves all metrics, and the decrease in test accuracy and PGD accuracy is not significant. Especially in the Fashion-MNIST dataset, after combining our method, the UNSATof the RS method in M3, $\varepsilon=0.3$ increases from 0.0 to 30%, which is a significant improvement considering the verification capability of the solver.
When combining our method with other training methods, the network’s verification performance is significantly improved; however, the test accuracy and robustness may be reduced. It needs a trade-off between verification performance and robustness depending on the application scenario when combining our method with other training methods.
[…]
M1 0.1 $1$ 97.61 100 3.24 3.24 51.16 89.8
$2^{r[i]}$ 98.61 100 3.22 3.22 52.18 85.9
$r[i]\cdot m[i]$ 98.88 98 3.05 3.12 53.66 75.3
C1 $\frac{2}{255}$ $1$ 68.10 36 23.56 42.40 95.46 66.55
$2^{r[i]}$ 69.53 35 35.83 63.57 97.46 62.05
$r[i]\cdot m[i]$ 71.15 32 44.53 77.45 97.83 51.37
Table 9: Networks trained with different $\gamma[i]$ on CIFAR-10 dataset and MNIST dataset.
[…]
M1 0.1 SABR 89.25 22 26.89 20.74 25.91 84.16
NBC 97.61 100 3.22 3.22 97.47 89.78
C1 $\frac{2}{255}$ SABR 59.57 50 6.55 3.49 48.53 87.57
NBC 62.48 54 6.69 5.02 51.55 79.60
Table 10: Comparison of networks trained with NBC and SABR on CIFAR-10 dataset and MNIST dataset.
Discussion for Hyperparameters
$\beta$ 0.0 1.0 3.0 6.0 9.0
Test Acc 76.0 72.2 70.0 67.1 66.0
UNSAT0 22.0 34.0 32.0 32.0
Stable4.8 42.1 60.9 71.8 74.3
Time (s) 44.4 45.4 39.9 23.7 19.4
[…]
$\text{PGD}_{100}$ 0.8 49.1 53.2 52.5 52.5
Table 11: Networks (C1 model) trained with NBC at $\varepsilon=2/255$ on CIFAR-10 dataset with different $\beta$.
Table 11 shows the results of C1 networks trained with NBC at $\varepsilon=2/255$ on CIFAR-10 dataset with different $\beta$. We evaluated the models with 50 verification tasks. Overall, as $\beta$ increases, the test accuracy decreases, UNSATincreases, stable neuron ratio increases, and verification time decreases, which is consistent with our expectations. The results show that a larger $\beta$ can improve the verification performance of the network, but it may reduce the accuracy of the network. Therefore, it is necessary to choose an appropriate $\beta$ according to the specific requirements of the application scenario.
Table 9 shows the results of networks trained with different $\gamma[i]$ on CIFAR-10 dataset and MNIST dataset. We evaluated the models with 50 verification tasks. As discussed in the main paper, neurons in smaller layers, when behavior is consistent, may constrain subsequent layers through propagation. Moreover, layers near the input and output often have fewer neurons, applying constraints to these layers can directly affect the forward or backward propagation process and accelerate the convergence of target loss. In contrast, middle layers typically have more neurons and are responsible for extracting complex features; over-constraining these layers could adversely affect the model’s expressive power. The results show that using a looser penalty on layers with more neurons may increase accuracy but decrease the proportion of stable neurons, which is consistent with our expectations. This indicates that it is necessary to choose an appropriate penalty term based on the scenario.
Compare with Certified Training
Table 10 shows the comparison of networks trained with NBC and SABR. We evaluated the models with 50 verification tasks. In the MINST dataset, the network trained with NBC achieves better performance in all metrics compared to the network trained with SABR. In the CIFAR-10 dataset, the network trained with NBC achieves better performance in terms of Test accuracy, UNSAT%, and accuracy under the PGD attack. These results show that under the same network architecture, our method—despite its lower computational cost and ease of implementation—achieves performance comparable to, or even surpasses, that of certified training methods which demand more computational resources.